Determine the wavelength in nm of light emitted by an electron as it falls from the 4th to the 2nd energy level.

Answer must be in computer format for scientific notation, e.g. 5.4E5


1
Expert's answer
2021-10-07T01:44:01-0400

The expression for the wavelength of radiation is, 

"\\frac{1}{\\lambda}= R(\\frac{1}{n^2_1}\n\n\n\n\u200b\u2212\\frac{1}{n^2_2}\n\n\u200b)" λ


Substitute values in the above expression,

"\\frac{1}{\u03bb}\n\n\n\n\u200b=109677\u00d7(\\frac{1}{2^2}\n\n\n\u200b\u2212\\frac{1}{4^2}\n\n\n\u200b)=109677\u00d7(\\frac{1}{4}-\\frac{\n1}{\n\u200b16}\n\n\n\u200b)"

"=109677\u00d7(\\frac{12}{64}\n\u200b)"



"\u03bb=4.86\u00d710^{\u22125}"


"cm=486\u00d710^{\u22129}\n."


"m=486nm"